A comparative balance sheet for Flounder Corporation is presented as follows. Assets Cash Accounts receivable Inventory Land1. 2. Net income for 2020 was $127,860. No gains or losses were recorded in 2020. Cash dividends of $63,790 were declared andFLOUNDER CORPORATION Statement of Cash Flows For the Year Ended December 31, 2020 Cash Flows from Operating Activities Net In$ Issued common stock to retire $ of bonds outstanding
Answer #1

Answer

Flounder corporation

Statement of cash flow

    For the Year Ended 31st December, 2020

Particulars $ $
Cash flow from operating activities
Net profit 127,860
Add depreciation 26,690
Add decrease in inventory 9,310
Less increase in account receivables (15,690)
Less decrease in account payable (13,310)
Cash flow from operating activities 134,860
Cash flow from investing activities
Purchase of equipment (59,690)
Sale of land 39,310
Cash flow from financing activities (20,380)
Cash flow from financing activities
Dividend paid (63,790)
Cash flow from financing activities (63,790)
Increase in cash 50,690
Beginning cash balance 22,000
Ending cash balance 72,690

Calculations

Increase in cash

= 134,860 + (20,380) + (63,790) = $50,690

Depreciation can be find by taking the difference between accumulated depreciation in 2019 and 2020.

Depreciation expense = 70,430 - 43,740 = $26,690

* Bond retirement by issuing common stock $50,000 is a non cash transaction, so it is not taken in cashflow statement.

Cash dividend is taken in financing activities as a outflow.

If YOU are like the answer please upvote.